OpenStudy (anonymous):

hen one U.S. Dollar is worth one point five two Australian Dollars, which of the following could be used as a conversion factor? nineteen one hundred twenty-fifths one and twelve twenty-fifths one and thirteen twenty-fifths five and one-fiftieth

OpenStudy (asnaseer):

1 Us Dollar = 1.52 Aus Dollars this means the conversion factor is 1.52 1.52 can also be written as:\[1.52=1+0.52=1\frac{52}{100}\]now just simplify the fraction 52/100